Some states have license plates formed by 3 letters followed by 3 digits. Both letters and the digits may be repeated.
If license plates are randomly assigned, what is the probability that a person would receive a license plate that contains all vowels and odd digits.

Answer by stanbon(75887)

Some states have license plates formed by 3 letters followed by 3 digits. Both letters and the digits may be repeated.
If license plates are randomly assigned, what is the probability that a person would receive a license plate that contains all vowels and odd digits.
---
Info:
5 vowels
5 odd digits
----------
# of ways to get 3 vowels: 5C3 = 5C2 = 10
# of ways to get 3 odd digits: 5C3 = 5C2 = 10
----
# of plates with all vowels and odd digits: 10*10 = 100
----
# of random plates: 26C3*10C3 = 312000
---
P(plate with vowels and odd digits) 100/312000 = 0.0003205...
Cheers,
Stan H.